OpenStudy (anonymous):

A line passes through the point 2,3 and has a slope of 4. Find the equation of the line. Answer Choices: y= 4x - 11 y= -4x - 5 y= -4x - 11 y = 4x - 5

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&\\ &(2\quad ,&3) \end{array} \\\quad \\ slope = m=4 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

whats answer?

OpenStudy (jdoe0001):

solve for "y"

OpenStudy (anonymous):

i dont understand this i would appreciate answer

OpenStudy (jdoe0001):

ever wonder why you weren't given the answer sheet along with the exercises?

OpenStudy (anonymous):

LOL i come on here for the answers not on how to do it f u

OpenStudy (jdoe0001):

the idea is that you understand it by solving it, not just transcribing things over

OpenStudy (anonymous):

is it b?